# Exact Modeling of Power Spectrum Multipole through Spherical Fourier-Bessel Basis

### v1

@@ Abstract
Conclusion and discussion– Through Eqs. 8 and 23, the main results of this letter, we have expressed the PSM and its covariance in terms of the generalized SFB power spectrum. The SFB basis naturally incorporates the WA effect present in the PSM without perturbative expansion of galaxy separation. It also provides a new way to do window convolution of the PSM through the SFB basis, without using the effective redshift approximation Castorina et al. (2019); Castorina and Di Dio (2022), which will potentially allow the use of large redshift bin(s) for cosmology with PSM.
@@
Since the mappings in Eqs. 8 and 23 are independent of cosmological models, we can add GR effects by including additional angular projection kernels $\Delta_{L}(k,z)$ Bonvin and Durrer (2011); Di Dio et al. (2013). In fact, the GR Doppler effect, which has the same form as Eq. 16 (just with different coefficients), has already been included in this work. The PSM-SFB mapping offers another non-perturbative method for modeling WA and GR effects on PSM in addition to the correlation function approach outlined in Ref. Castorina and Di Dio (2022).
@@
We are now able to validate the accuracy of the numerical codes for GR effects by comparing results from two different approaches, which will be important for upcoming LSS surveys with the aim to measure $\sigma(f_{\mathrm{NL}})\sim O(1)$ Foglieni et al. (2023). The natural next step for our work will be to add the remaining GR terms to the gSFB power spectrum, which will then enable the computation of the PSM covariance with GR effects, and allow us to forecast the measurability of GR effects and assess their impact on $f_{\mathrm{NL}}$ constraints in the Fourier space for future surveys.
@@
Our result Eq. 8 suggests that PSM serves as a natural compression of the SFB power spectrum. It will be interesting to study the loss of information caused by such compression. Yamamoto-like estimators assume a pair LOS, which causes information loss for galaxies with large angular separations. In contrast, SFB allows for a different LOS for every galaxy and fully preserves the redshift evolution, serving as a more optimal statistic. This motivates a search for better compression schemes of the SFB on large scales to potentially improve upon the PSM.
@@
To apply our formalism for modeling the PSM on realistic data, we have to account for realistic window and efficiently evaluate the generalized SFB power spectrum. We have so far restricted ourselves to separable windows. For non-separable windows where the redshift distribution is anisotropic, we will need to extend Eqs. 13 and 23 and evaluate the window convolution of the SFB mode under the generic case as suggested in Ref. Grasshorn Gebhardt and Doré (2021).
@@
Our mapping Eq. 8 considered the PSM and the gSFB as theoretical quantities. However, it also applies to them as estimators used on data. It will be interesting to explore the use of the mapping at the estimator level to validate the consistency of the measurements of two statistics. As first suggested in Ref. Castorina and White (2018), Eq. 8 can be potentially used to control various systematic effects. In the SFB basis, there is a clear separation between radial and angular scales, so that we can remove any systematic-dominated angular modes from the Yamamoto estimator by removing a few SFB modes.
@@
Furthermore, our work can be extended to the bispectrum. The numerical evaluation of the SFB bispectrum has been recently achieved in Ref. Benabou et al. (2023), which will allow for an accurate modeling of WA, GR and window effects present in the bispectrum multipoles on large scales without taking the perturbative approach typically used in literature Pardede et al. (2022, 2023); Noorikuhani and Scoccimarro (2023).
@@
Acknowledgements– We thank the SPHEREx team for useful discussion. Part of this work was done at Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004).
@@
Appendix– Here we provide a derivation of the PSM-SFB mapping in Eq. 8. We first expand all the exponential terms in Eq. 5 with
@@
Applying the gSFB decomposition of Eq. 2 to the windowed field $\delta^{\rm W}(\mathbb{x})$ and using the windowed gSFB (pseudo) power spectrum in Eq. 4, we have
@@
$\frac{I_{22}}{(2L+1)}P_{L}(k)$ $=\sum_{a,b}i^{-a+b}(2a+1)(2b+1)\int_{\mathbb{x},\mathbb{x}^{\prime}}j_{a}(kx)j_{b}(kx^{\prime})\langle\delta^{{\rm W}}(\mathbb{x})\delta^{{\rm W}}(\mathbb{x}^{\prime})\rangle\int_{\hat{\mathbf{k}}}\mathcal{L}_{a}(\hat{\mathbf{k}}\cdot\hat{\mathbb{x}})\mathcal{L}_{b}(\hat{\mathbf{k}}\cdot\hat{\mathbb{x}}^{\prime})\mathcal{L}_{L}(\hat{\mathbf{k}}\cdot\hat{\mathbb{x}})$ (31)
$=(4\pi)^{2}\sum_{a,b}i^{-a+b}(2a+1)\begin{pmatrix}a&L&b\\ 0&0&0\end{pmatrix}^{2}\sum_{m_{b}}\langle\int_{\mathbb{x}}j_{a}(kx)\delta^{{\rm W}}(\mathbb{x})Y_{bm_{b}}^{*}(\hat{\mathbb{x}})\int_{\mathbb{x}^{\prime}}j_{b}(kx^{\prime})\delta^{{\rm W}}(\mathbb{x}^{\prime})Y_{bm_{b}}(\hat{\mathbb{x}}^{\prime})\rangle\,,$ (34)
which then becomes Eq. 8.

### v2

@@ V Conclusion and Discussion
Through Eqs. 6 and 16, the main results of this work, we have expressed the PSM and its covariance in terms of the generalized SFB power spectrum. The SFB basis naturally incorporates the WA effect present in the PSM without perturbative expansion of galaxy separation. It also provides a new way to do window convolution of the PSM through the SFB basis, without using the effective redshift approximation Castorina et al. (2019); Castorina and Di Dio (2022), which will potentially allow the use of large redshift bin(s) for cosmology with PSM.
@@
Since the mappings in Eqs. 6 and 16 are independent of cosmological models, we can add GR effects by including additional angular kernels $\Delta_{L}(k,z)$ Bonvin and Durrer (2011); Di Dio et al. (2013). In fact, the GR Doppler effect, which has the same form as Eq. 14 (just with different coefficients), has already been included in this work. The SFB-to-PSM mapping offers another non-perturbative method for modeling WA and GR effects on PSM in addition to the correlation function approach outlined in Ref. Castorina and Di Dio (2022).
@@
We are now able to validate the accuracy of the numerical codes for GR effects by comparing results from two different approaches, which will be important for upcoming LSS surveys with the aim to measure $\sigma(f_{\mathrm{NL}})\sim O(1)$ Foglieni et al. (2023). The natural next step for our work will be to add the remaining GR terms to the gSFB power spectrum, which will enable the computation of the PSM covariance with GR effects, and allow us to forecast the measurability of GR effects and assess their impact on $f_{\mathrm{NL}}$ constraints in the Fourier space for future surveys.
@@
Our result Eq. 6 suggests that PSM serves as a natural compression of the SFB power spectrum. It will be interesting to study the loss of information caused by such compression. Yamamoto-like estimators assume a pair LOS, which can cause information loss for galaxies with large angular separations. In contrast, SFB allows for a different LOS for every galaxy and fully preserves the redshift evolution, serving as a more optimal statistic. This motivates a search for better compression schemes of the SFB on large scales to potentially improve upon the PSM.
Alternatively, to avoid the loss of information and the non-linear modeling of SFB power spectrum, one can use a hybrid estimator as suggested in Ref. Wang et al. (2020), that is using the SFB basis on large scales and the Cartesian basis on smaller scales. Employing the hybrid method on upcoming LSS surveys will require further development in the approach, and our exact modeling of PSM at large scales offers the baseline. The SFB-to-PSM mapping can help model the covariance of the hybrid estimator, establish consistency between the two basis, and optimize the transition scale between the two basis to balance constraining power and computational efficiency. Using PSM on large scales still has the advantage of reducing the number of data vectors and allowing the use of a single statistic and a single estimator (the Yamamoto estimator) across all scales, along with simpler joint modeling with bispectrum multipoles.
@@
To apply our formalism for modeling the PSM on realistic data, we have to account for realistic window and efficiently evaluate the generalized SFB power spectrum. We have so far restricted ourselves to separable windows. For non-separable windows where the redshift distribution is anisotropic, we will need to extend Eqs. 11 and 16 and evaluate the window convolution of the SFB mode under the generic case as considered in Ref. Grasshorn Gebhardt and Doré (2021). Non-separable windows will cause mixing of both angular and radial modes, which will be significantly more expensive to evaluate. However, the evaluation of the generic window convolution is potentially achievable for the largest scales where there are only a few SFB modes and the effect of a non-separable window is the most prominent.
Throughout the work, we have only considered the linear modeling of PSM. For even multipoles, the wide-angle effects are only important at the largest scales where the linear-order theory is sufficient, so one can use the SFB-to-PSM mapping developed in this work at the largest scales for WA modeling and switch to the standard plane-parallel approximation at the quasi-linear scales for 1-loop calculations. In contrast, the odd multipoles vanish under the plane-parallel approximation for auto-correlation and have to be modeled under the wide-angle regime at all scales. One can then adopt the perturbative WA approach, which only breaks down at the largest scales as shown in Ref. Benabou et al. (2024), at the quasi-linear scales for 1-loop calculations Noorikuhani and Scoccimarro (2023). Alternatively, it might be interesting to explore the challenging task of evaluating the SFB power spectrum at the 1-loop order, which can then be mapped to PSM for obtaining both odd and even multipoles beyond the linear theory.
@@
Our mapping Eq. 6 considered the PSM and the gSFB as theoretical quantities. However, it also applies to them as estimators used on data. It will be interesting to explore the use of the mapping at the estimator level to validate the consistency of the measurements of two statistics. As suggested in Ref. Castorina and White (2018), Eq. 6 can be potentially used to control various systematic effects. In the SFB basis, there is a clear separation between radial and angular scales, so we can remove any systematic-dominated angular modes from the Yamamoto estimator by removing a few SFB modes.
@@
Furthermore, our work can be extended to the bispectrum. The numerical evaluation of the SFB bispectrum has been recently achieved in Ref. Benabou et al. (2023), which will allow for an accurate modeling of WA, GR and window effects present in the bispectrum multipoles on large scales without taking the perturbative approach used in literature Pardede et al. (2022, 2023); Noorikuhani and Scoccimarro (2023).
Acknowledgements
@@
We thank Joshua Benabou, Isabel Sands, Richard Feder, Mike S. Wang, and the SPHEREx cosmology team for useful discussion and feedback. We acknowledge support from the SPHEREx project under a contract from the NASA/GODDARD Space Flight Center to the California Institute of Technology. Part of this work was done at Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (Contract No. 80NM0018D0004).
Appendix A Derivation of SFB-to-PSM Mapping
@@
Here we provide a derivation of the SFB-to-PSM mapping in Eq. 6. We first expand all the exponential terms in Eq. 5 with
@@
This gives us:
@@
$\frac{I_{22}}{(2L+1)}P_{L}(k)$ $=\sum_{a,b}i^{-a+b}(2a+1)(2b+1)\int_{\mathbb{x},\mathbb{x}^{\prime}}j_{a}(kx)j_{b}(kx^{\prime})\langle\delta^{{\rm W}}(\mathbb{x})\delta^{{\rm W}}(\mathbb{x}^{\prime})\rangle\int_{\hat{\mathbf{k}}}\mathcal{L}_{a}(\hat{\mathbf{k}}\cdot\hat{\mathbb{x}})\mathcal{L}_{b}(\hat{\mathbf{k}}\cdot\hat{\mathbb{x}}^{\prime})\mathcal{L}_{L}(\hat{\mathbf{k}}\cdot\hat{\mathbb{x}})$
$=(4\pi)^{2}\sum_{a,b}i^{-a+b}(2a+1)\begin{pmatrix}a&L&b\\ 0&0&0\end{pmatrix}^{2}\sum_{m_{b}}\langle\int_{\mathbb{x}}j_{a}(kx)\delta^{{\rm W}}(\mathbb{x})Y_{bm_{b}}^{*}(\hat{\mathbb{x}})\int_{\mathbb{x}^{\prime}}j_{b}(kx^{\prime})\delta^{{\rm W}}(\mathbb{x}^{\prime})Y_{bm_{b}}(\hat{\mathbb{x}}^{\prime})\rangle\,.$ (22)
Applying the gSFB decomposition of Eq. 2 to the windowed field $\delta^{\rm W}(\mathbb{x})$ and using the windowed gSFB (pseudo) power spectrum defined in Eq. 4, Eq. 22 will simplify to Eq. 6.
Appendix B Derivation of PSM Covariance
Here we give a derivation of the PSM Gaussian covariance in Eq. 16. Starting with the covariance definition in Eq. 15, we split the 4-point function assuming Gaussianity and substitute $F_{L}(k)$ with Eq. 13:
$\frac{I_{22}^{2}}{(2L_{1}+1)(2L_{2}+1)}\textbf{C}^{\textup{G}}_{L_{1}L_{2}}(k_{1},k_{2})=\int_{\hat{\mathbf{k}}_{1},\hat{\mathbf{k}}_{2}}\bigg{[}\langle F_{L_{1}}(\mathbf{k}_{1})F_{0}(-\mathbf{k}_{2})\rangle\langle F_{L_{2}}(\mathbf{k}_{2})F_{0}(-\mathbf{k}_{1})\rangle$
$\qquad+\langle F_{L_{1}}(\mathbf{k}_{1})F_{L_{2}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{k}_{2}})\rangle\langle F_{0}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\mathbf{k}_{2}})F_{0}(-\mathbf{k}_{1})\rangle\bigg{]}$
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\hat{\mathbf{k}}_{1},\hat{\mathbf{k}}_{2}}\int_{\mathbb{x}_{1},\mathbb{x}^{\prime}_{1},\mathbb{x}_{2},\mathbb{x}^{\prime}_{2}}e^{-i\mathbf{k}_{1}\cdot(\mathbb{x}_{1}-\mathbb{x}^{\prime}_{1})}\langle\delta^{\rm W}(\mathbb{x}_{1})\delta^{\rm W}(\mathbb{x}^{\prime}_{2})\rangle\langle\delta^{\rm W}(\mathbb{x}^{\prime}_{1})\delta^{\rm W}(\mathbb{x}_{2})\rangle\mathcal{L}_{L_{1}}(\hat{\mathbb{x}}_{1}\cdot\hat{\mathbf{k}}_{1})}$
$\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Big{[}e^{-i\mathbf{k}_{2}\cdot(\mathbb{x}_{2}-\mathbb{x}^{\prime}_{2})}\mathcal{L}_{L_{2}}(\hat{\mathbb{x}}_{2}\cdot\hat{\mathbf{k}}_{2})+e^{i\mathbf{k}_{2}\cdot(\mathbb{x}_{2}-\mathbb{x}^{\prime}_{2})}\mathcal{L}_{L_{2}}(\hat{\mathbb{x}}^{\prime}_{2}\cdot\hat{\mathbf{k}}_{2})\Big{]}}\,.$ (23)
We consider the first term in the above expression (the $\mathcal{L}_{L_{1}}(\hat{\mathbb{x}}_{1}\cdot\hat{\mathbf{k}}_{1})\mathcal{L}_{L_{2}}(\hat{\mathbb{x}}_{2}\cdot\hat{\mathbf{k}}_{2})$ term, denoted as $\textbf{C}^{\textup{G-1}}_{L_{1}L_{2}}(k_{1},k_{2})$). Expanding all the exponentials with the plane wave expansion Eq. 20 and using Eq. 21 to evaluate the $\hat{\mathbf{k}}_{1}$ and $\hat{\mathbf{k}}_{2}$ integrals, we obtain
$\frac{I_{22}^{2}}{(2L_{1}+1)(2L_{2}+1)}\textbf{C}^{\textup{G-1}}_{L_{1}L_{2}}(k_{1},k_{2})=(4\pi)^{4}\sum_{a,b,c,d,L_{1},L_{2}}i^{-a+b-c+d}(2a+1)(2c+1)\begin{pmatrix}a&L_{1}&b\\ 0&0&0\end{pmatrix}^{2}\begin{pmatrix}c&L_{2}&d\\ 0&0&0\end{pmatrix}^{2}\sum_{m_{b}.m_{d}}\int_{\mathbb{x}_{1},\mathbb{x}^{\prime}_{2}}$
$\quad\langle\delta^{\rm W}(\mathbb{x}_{1})\delta^{\rm W}(\mathbb{x}^{\prime}_{2})\rangle j_{a}(k_{1}x_{1})j_{d}(k_{2}x_{2}^{\prime})Y^{*}_{bm_{b}}(\hat{x}_{1})Y_{dm_{d}}(\hat{x}^{\prime}_{2})\int_{\mathbb{x}_{1}^{\prime},\mathbb{x}_{2}}\langle\delta^{\rm W}(\mathbb{x}^{\prime}_{1})\delta^{\rm W}(\mathbb{x}_{2})\rangle j_{b}(k_{1}x^{\prime}_{1})j_{c}(k_{2}x_{2})Y_{bm_{b}}(\hat{x}^{\prime}_{1})Y^{*}_{dm_{d}}(\hat{x}_{2})\,.$
We can apply the same procedure to the second term containing $\mathcal{L}_{L_{1}}(\hat{\mathbb{x}}_{1}\cdot\hat{\mathbf{k}}_{1})\mathcal{L}_{L_{2}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mathbb{x}}^{\prime}_{2}\cdot\hat{\mathbf{k}}_{2}})$ in Eq. 23. We then apply the gSFB decomposition of Eq. 2 to the windowed field $\delta^{\rm W}(\mathbb{x})$ and the Gaussian covariance becomes:
$\frac{I_{22}^{2}}{(2L_{1}+1)(2L_{2}+1)}\textbf{C}^{\textup{G}}_{L_{1}L_{2}}(k_{1},k_{2})=(4\pi)^{4}\sum_{a,b,c,d,L_{1},L_{2}}i^{-a+b-c+d}(2a+1)\begin{pmatrix}a&L_{1}&b\\ 0&0&0\end{pmatrix}^{2}\begin{pmatrix}c&L_{2}&d\\ 0&0&0\end{pmatrix}^{2}\Bigg{[}(2c+1)\sum_{m_{b},m_{d}}$
$\quad\langle\delta_{bm_{b}}^{a,{\rm W}}(k_{1})\delta_{dm_{d}}^{d*,{\rm W}}(k_{2})\rangle\langle\delta_{bm_{b}}^{b*,{\rm W}}(k_{1})\delta_{dm_{d}}^{c,{\rm W}}(k_{2})\rangle+(2d+1){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(-1)^{c+d}}\sum_{m_{b},m_{c}}\langle\delta_{bm_{b}}^{a,{\rm W}}(k_{1})\delta_{cm_{c}}^{d*,{\rm W}}(k_{2})\rangle\langle\delta_{bm_{b}}^{b*,{\rm W}}(k_{1})\delta_{cm_{c}}^{c,{\rm W}}(k_{2})\rangle\Bigg{]}\,,$ (24)
where we have expressed the PSM Gaussian covariance in terms of the gSFB modes.
We now assume the separability of the window function into the angular and radial parts $W(\mathbb{x})=R(x)M(\hat{\mathbb{x}})$. We consider the impact of the angular mask on the gSFB mode, and we find
$\delta_{\ell m}^{a,{\rm W}}(k)=\int_{x}j_{L}(kx)\int_{\hat{\mathbb{x}}}Y^{*}_{\ell m}(\hat{\mathbb{x}})M(\hat{\mathbb{x}})\sum_{\ell^{\prime}m^{\prime}}Y_{\ell^{\prime}m^{\prime}}(\hat{\mathbb{x}})\delta^{\rm R}_{\ell^{\prime}m^{\prime}}(x)=\sum_{\ell^{\prime}m^{\prime}}M^{\ell m}_{\ell^{\prime}m^{\prime}}\delta_{\ell^{\prime}m^{\prime}}^{a,{\rm R}}(k),$ (25)
where the superscript R indicates the presence of only radial selection without any angular mask, and $M^{\ell m}_{\ell^{\prime}m^{\prime}}$ is the double spherical harmonic transform of the angular mask defined in Eq. 18. Therefore,
$\langle\delta_{bm_{b}}^{x,{\rm W}}(k_{1})\delta_{dm_{d}}^{y*,{\rm W}}(k_{2})\rangle=\sum_{\ell,m_{\ell}}M_{\ell m_{\ell}}^{bm_{b}}\sum_{u,m_{u}}M^{um_{u}}_{dm_{d}}\langle\delta_{\ell m_{\ell}}^{x,{\rm W}}(k_{1})\delta_{um_{u}}^{y*,{\rm W}}(k_{2})\rangle=\sum_{\ell,m_{\ell}}M_{\ell m_{\ell}}^{bm_{b}}M^{\ell m_{\ell}}_{dm_{d}}C_{\ell}^{xy,{\rm R}}(k_{1},k_{2})\,,$ (26)
where we have used Eq. 3 that defines the gSFB power spectrum under azimuthal symmetry.
We now use the above Eq. 26 to simplify the following expression contained in the first term of Eq. 24:
$\sum_{m_{b},m_{d}}\langle\delta_{bm_{b}}^{a,{\rm W}}(k_{1})\delta_{dm_{d}}^{d*,{\rm W}}(k_{2})\rangle\langle\delta_{bm_{b}}^{b*,{\rm W}}(k_{1})\delta_{dm_{d}}^{c,{\rm W}}(k_{2})\rangle$
$=$ $\sum_{\ell_{1},\ell_{2},m_{b},m_{d},m_{1},m_{2}}M_{\ell_{1}m_{1}}^{bm_{b}}M^{\ell_{1}m_{1}}_{dm_{d}}C_{\ell_{1}}^{ad,{\rm R}}(k_{1},k_{2})M^{\ell_{2}m_{2}}_{bm_{b}}M_{\ell_{2}m_{2}}^{dm_{d}}C_{\ell_{2}}^{bc,{\rm R}}(k_{1},k_{2})$
$=$ $\sum_{\ell_{1},\ell_{2}}S_{b\ell_{1}d\ell_{2}}C_{\ell_{1}}^{ad,{\rm R}}(k_{1},k_{2})C_{\ell_{2}}^{bc,{\rm R}}(k_{1},k_{2})\,,$ (27)
where $S_{b\ell_{1}d\ell_{2}}$ is the chain of window defined in Eq. 17. One can similarly simplify the second term of Eq. 24, and the covariance expression in Eq. 24 will reduce to Eq. 16 under any separable window.